### v1

# Efficient Data Collection for Robotic Manipulation via Compositional Generalization

Abstract
Data collection has become an increasingly important problem in robotic manipulation, yet there still lacks much understanding of how to effectively collect data to facilitate broad generalization. Recent works on large-scale robotic data collection typically vary a wide range of environmental factors during data collection, such as object types and table textures. While these works attempt to cover a diverse variety of scenarios, they do not explicitly account for the possible compositional abilities of policies trained on the data. If robot policies are able to compose different environmental factors of variation (e.g., object types, table heights) from their training data to succeed when encountering unseen factor combinations, then we can exploit this to avoid collecting data for situations that composition would address. To investigate this possibility, we conduct thorough empirical studies both in simulation and on a real robot that compare data collection strategies and assess whether visual imitation learning policies can compose environmental factors. We find that policies do exhibit composition, although leveraging prior robotic datasets is critical for this on a real robot. We use these insights to provide better practices for in-domain data collection by proposing data collection strategies that exploit composition, which can induce better generalization than naïve approaches for the same amount of effort during data collection. We further demonstrate that a real robot policy trained on data from such a strategy achieves a success rate of 77.5% when transferred to entirely new environments that encompass unseen combinations of environmental factors, whereas policies trained using data collected without accounting for environmental variation fail to transfer effectively, with a success rate of only 2.5%. We provide videos at our project website. [^1] http://iliad.stanford.edu/robot-data-comp/
I Introduction
For robots to be practical and useful, they must robustly handle the wide variety of conditions they will encounter in the world. Recent advances in machine learning have shown that leveraging diverse, internet-scale data can be very effective in facilitating broad generalization [7, 38], as such data captures much of the complexity and variation in the domain that models will need to reason about. It therefore has been of great interest in the robotics community to apply a similar recipe to robotics, specifically using end-to-end imitation learning, which holds promise to effectively scale with large, diverse datasets to achieve broad generalization. However, we lack access to existing internet-scale robotics data, and furthermore, robotics data – especially human-collected demonstrations – is often limited and challenging to collect. While recent works have made significant progress in scaling real-world robotic manipulation datasets [44, 17, 35, 2], their scale is still much less than the datasets typically used for pre-training in computer vision and natural language processing. As a result, policies trained on such datasets still exhibit unsatisfactory zero-shot performance when deployed in new environments, often necessitating additional in-domain data collection for generalization. Given the lack of large-scale robotics data and the high cost of data collection, it is essential to focus on how to best collect this in-domain data.
Figure 1: We investigate if robot policies can compose environmental factors of variation (e.g., object types, table heights) in their in-domain training data (1, 2), and if using prior data can be helpful for enforcing such generalization. We propose efficient in-domain data collection strategies guided by the ability of policies to reason about unseen combinations of factor values (3).
When collecting robotics data to account for a given range of scenarios, it is often infeasible to cover all possible settings. Instead, it would be desirable to exploit the compositional generalization capabilities of the policies being trained, i.e., their ability to reason effectively about unseen combinations of environmental factors of variation. For example, for the task of picking up different objects from tables of different heights, as shown in Fig. 1, it would simply not be scalable for humans to collect data for all combinations of desired objects and table heights. Instead, it would be much more feasible if one could prioritize collecting data to cover all individual examples of objects and table heights (1 and 2 in Fig. 1), but not necessarily all of their combinations, and then have a policy trained on this data generalize to unseen combinations (3 in Fig. 1). By doing so, we could drastically reduce the amount of data needed.
While prior work has studied generalization to unseen environmental factors in robotics [47, 46], there has not been as much focus on understanding compositional generalization in robotics, particularly for end-to-end imitation learning. Compositional generalization has been previously studied in other machine learning domains [19, 1, 25, 22, 23, 41, 4, 49], with results often suggesting that end-to-end neural models can struggle to achieve this. However, recent work has shown that large language models can exhibit strong compositional abilities, likely due to the diversity and scale of their training data [51, 13]. Ideally, prior datasets in robotics could also facilitate composition for robotics, but we still lack much understanding as to where these datasets benefit generalization.
Our key idea is that there is more guidance that can be provided for efficient data collection in robotics. While there has been a recent trend in scaling up robotic data collection, most prior works attempt to simply collect diverse data covering all possible combinations of different factors of variation [44, 17, 35, 2]. However, we hypothesize that imitation learning policies – similar to their large language model counterparts – can potentially exhibit compositional generalization, in particular when leveraging prior robotics datasets. In this work, we investigate this hypothesis, with the goal of prescribing better practices for robotic data collection.
Through extensive simulated and real experiments, we study the proficiency of visual imitation learning policies when evaluated on unseen combinations of environmental factors. We consider a wide variety of factors that are broadly relevant in robotic manipulation and have been previously studied, including object position, object type, table texture, table height, camera position, and distractor objects. We find significant evidence of compositional generalization with imitation policies when varying these factors. Given this evidence, we propose in-domain data collection strategies that exploit this generalization – providing more systematic guidelines that can help reduce data collection burden for roboticists. On a real robot, we find that a policy using data from such a strategy succeeds in 59/90 settings that assess composition, compared to 22/90 when using the same amount of in-domain data, but without explicit variation. We also find that using prior robot data – in our case, BridgeData V2 [44] – is critical for this composition, with performance dropping to 28/90 without it. Furthermore, we evaluate policies when transferred to entirely new environments that encompass unseen combinations of environmental factors, and find that our best policy achieves a success rate of 77.5%, compared to 27.5% when there is no prior data, and 2.5% when there is no variation in the in-domain data from our training environments.
II Related Work
In this section, we review prior work on robotic data collection, generalization in robot learning, and compositional generalization more broadly in machine learning.
Data Collection in Robot Learning. Prior works have studied data collection methods in robot learning to improve generalization. Much of this work focuses on scaling up robot data collection [37, 15, 27, 32, 42, 10, 16, 5, 44, 17, 2]. Results from these works have demonstrated that data with greater object and task diversity can improve policy robustness, suggesting that data should be collected with this diversity in mind. In our work, we study how using such datasets as prior data affects compositional generalization. However, results from these works also show that policies trained on these datasets often exhibit poor zero-shot performance in new scenarios, suggesting that in-domain data collection is still often needed in practice. Unlike these prior works, our work focuses on how to collect this in-domain data, by studying how in-domain data composition affects generalization at a more nuanced level, and systematically considering a wide variety of specific environmental factors of variation.
Other works on data collection in imitation learning focus on expanding state coverage via interactive imitation learning [40, 26, 21], or more recently improving other notions of data quality [3]. These works mainly consider the impact of data on the distribution shift problem [40], to improve policy robustness to the state distribution the agent will encounter online. However, distribution shift in robotics can also occur due to environmental variations. We instead focus on addressing this form of distribution shift, through offline data collection with active variation of environmental factors.
Generalization in Robot Learning. Much prior work has studied improving generalization in robot learning, such as by using pre-trained visual representations [36, 33, 30, 39, 11, 31, 20], or pre-training on diverse robot data [35, 34]. In our work, we focus on assessing compositional generalization of existing imitation learning methods. Some prior works have similarly studied the robustness of visual imitation learning models to different environmental factor shifts [47, 46, 8]. However, these works only consider generalization to out-of-distribution values for one or two factors at a time. Our work instead assesses compositional generalization, where policies must reason about unseen combinations of factor values.
Compositional Generalization. Compositional generalization has been studied extensively in many machine learning areas, including visual reasoning [41, 49], image generation [14, 28], natural language processing [25, 22, 23], and visual question answering [19, 1, 4]. Much of these works demonstrate that end-to-end neural models often struggle with composition, although sometimes they can exhibit better composition than methods specifically intended to facilitate it, such as unsupervised representation learning [41]. However, more recent work has shown that large language models can possess strong compositional abilities, likely owing to their internet-scale pre-training data [51, 13]. In our work, we study if leveraging prior data can also be beneficial for composition in robotics.
Compositional generalization has also been investigated in robotics. This includes work on modular approaches for promoting composition [12, 48, 24, 45], and end-to-end architectures with inductive biases that benefit composition [50]. However, these works primarily study composition at the semantic level, for understanding concepts such as text instructions, high-level skills, or object properties and relationships. Unlike these works, we study composition of environmental factors, which may significantly change visual observations, or the low-level physical motions needed for a task. In addition, we study composition for existing end-to-end imitation learning methods. Such approaches have become popular in robotics and hold promise to scale effectively with data.
There has been some evidence of compositional generalization in end-to-end robotic imitation learning models. This includes composition of tasks in prior data with the conditions of a target domain [16], objects with manipulation skills [5, 43], semantic concepts from internet data with manipulation skills [6], and behaviors and semantic concepts across robot embodiments [5, 35]. However, our work provides more systematic and fine-grained analysis on when composition is possible, and considers a greater variety of environmental factors. Furthermore, unlike these works, we focus on using this analysis to guide better practices for data collection that take composition into account. We believe these differences allow our insights to be leveraged for more effective data collection in real-world settings.
III Exploiting Compositional Generalization for Efficient Data Collection
III-A Problem Statement
We consider the goal-conditioned imitation learning setting, where the objective is for a robot to reach a goal. We formalize this by defining the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},H,\mathcal{F}^{N},p(f),\rho(s_{0}|f),\mu(\cdot|f))$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $\mathcal{T}(s^{\prime}|s,a)$ is the transition dynamics function, and $H$ is the horizon length. We refer to $\mathcal{F}^{N}\subset\mathbb{Z}^{N}$ as the factor space, which captures changes in the environment along $N$ different axes that the robot could encounter, which we refer to as factors. For example, the first component of $\mathcal{F}^{N}$ could capture variation in object type, and the second could capture variation in table height, etc. For simplicity, we assume each component of $\mathcal{F}^{N}$ consists of $k$ possible discrete values, e.g., there are $k$ possible object types the robot may encounter. We refer to these as factor values. Therefore, each element $f\in\mathcal{F}^{N}$ defines a combination of factor values for the environment, e.g., a specific object type and table height the robot may encounter. $p(f)$ is the distribution of factor value combinations, and $\rho(s_{0}|f)$ is the initial state distribution, conditioned on a combination of factor values. We adopt this formalism to express that factor values in the environment can affect tasks by inducing different initial states. Similarly, $\mu(f):\mathcal{F}^{N}\rightarrow S$ is a deterministic function that maps a combination of factor values to the desired goal state $g\in S$, as the factor values can determine what the goal state is for a task. We aim to learn a goal-conditioned policy $\pi(\cdot|s,g):S\times S\rightarrow\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ is the probability simplex over the action space. We aim to maximize the probability of reaching the desired goal, under the distribution of possible factor values:
$J(\pi)=\mathbb{E}_{f\sim p(f)}[P_{\pi}(s_{H}=\mu(f))],$ (1)
where $P_{\pi}(s_{H}=\mu(f))$ is the probability that the state visited by policy $\pi$ at the final timestep $H$ is the goal state $g=\mu(f)$.
To learn the policy $\pi$, we assume a dataset of $M$ expert demonstrations $\mathcal{D}_{M}=(\tau_{1},\dots,\tau_{M})$. Each demonstration $\tau_{i}=\{(s_{1},a_{1}),\dots,(s_{T_{i}},a_{T_{i}})\}$ is a sequence of state-action pairs of length $T_{i}$, produced by sampling actions from an expert policy $\pi_{E}(\cdot|s,g)$ through dynamics $\mathcal{T}(s^{\prime}|s,a)$. Each demonstration also has an associated combination of factor values $f_{i}$, which defines both the initial state distribution the demonstration was collected from, and the goal state $g=\mu(f)$ the expert policy aimed to reach. We can then use this dataset to learn a policy $\pi$ using goal-conditioned behavior cloning [29].
III-B Objective
We aim to train robotic manipulation policies using imitation learning that will robustly handle combinations of factor values from some desired distribution $p(f)$, by collecting demonstrations with associated factor values $f$ according to some strategy. We assume without loss of generality that $p(f)=\text{Uniform}(\mathcal{F}^{N})$, such that we desire to handle all possible combinations of factor values in $\mathcal{F}^{N}$. The most direct way to achieve this robustness is sufficiently covering all of $\mathcal{F}^{N}$, such as by sampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ enough times, or even by collecting demonstrations for all $f\in\mathcal{F}^{N}$. However, as the cardinality of $\mathcal{F}^{N}$ increases exponentially with the number of factors $N$, collecting demonstrations for all possible factor value combinations is often impractical. Furthermore, constantly resampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ can often be expensive or challenging in practice, as each factor change often requires some degree of manual effort.
Rather than attempting to collect data for all combinations of factor values in $\mathcal{F}^{N}$, it would be desirable to instead collect data with a focus on simply capturing all individual factor values, and then exploit compositional generalization of the learned policy to perform well on unseen combinations. With $N$ different factors and $k$ possible values each, collecting data covering all combinations would require $\mathcal{O}(k^{N})$ factor changes in the environment, while covering all individual factor values would only require $\mathcal{O}(kN)$, a dramatic reduction. However, this is a sensible option only if compositional generalization is actually achieved by the learned policy. As prior work has shown mixed results on the compositional abilities of end-to-end neural models, it remains important to investigate when compositional generalization happens in imitation learning for robotic manipulation. Therefore, to exploit compositional generalization for more efficient data collection, we must answer the following questions:
  1. 1.
    When do robotic imitation learning policies exhibit compositional generalization?
  2. 2.
    What are effective data collection strategies to exploit composition, such that we can achieve broad generalization while reducing data collection effort?
We answer these questions by conducting extensive experiments in both simulation and on a real robot, investigating the compositional abilities of imitation learning policies for a variety of factors. We compare data collection strategies intended to exploit composition by optimizing for coverage of individual factor values, as opposed to naïvely optimizing for coverage of all factor combinations.
Figure 2: Visualization of our data collection strategies with $N=2$ factors. Each axis consists of possible values for a factor. Each green dot indicates that the strategy captures a specific combination of factor values represented by it, and each pink dot represents a combination that compositional generalization may address. We name our strategies based on the patterns in this visualization.
III-C Data Collection Strategies
We describe the data collection strategies that we consider. We visualize these strategies for when the number of factors $N=2$ in Fig. 2, but these can easily be extended for $N>2$. The strategies involve periodically setting factor values in the environment during data collection. To quantify effort, we consider the total number of factor changes made in the environment. Note that changing multiple factors at a time involves multiple changes, e.g., setting the environment to have different values for all factors involves $N$ changes.
  1. 1.
    Complete: This strategy covers all combinations of factor values in $\mathcal{F}^{N}$. As this requires $\mathcal{O}(k^{N})$ factor changes, this is often deemed infeasible in practice.
  2. 2.
    Random: This strategy periodically resamples an entire random combination of factor values $f$ from all possible $f\in\mathcal{F}^{N}$ without replacement. This strategy will eventually cover all of $\mathcal{F}^{N}$, but may achieve generalization inefficiently by not actively leveraging composition.
  3. 3.
    Single Factor: This strategy only varies one factor during data collection, while keeping the values of all other factors the same as some base factor values $f^{*}$.
  4. 4.
    Diagonal: This strategy resamples entire combinations of factor values $f$, but only samples $f$ where each factor value has never been seen. This requires $\mathcal{O}(kN)$ factor changes to cover all values, the fewest possible.
  5. 5.
    L: This strategy varies only one factor at a time. Starting from base factor values $f^{*}$, the possible combinations of factor values $f$ only deviate from $f^{*}$ by one factor. This also covers all factor values with $\mathcal{O}(kN)$ factor changes.
  6. 6.
    Stair: This strategy starts at $f^{*}$, and then cyclically varies one factor at a time, while preserving the values for all other factors. This also covers all factor values with $\mathcal{O}(kN)$ changes. However, compared to Diagonal, it covers more combinations of factor values, and compared to L, it captures more diverse combinations.
  7. 7.
    No Variation: This baseline involves only collecting data for a single base combination of factor values $f^{*}$.
Stair, L, and Diagonal are intended to exploit compositional generalization by prioritizing covering all individual factor values with the fewest amount of factor changes. We note that using factor changes as a measure of effort is a rough approximation. For example, this may not be accurate when it is more challenging to vary two factors together than separately. This could be the case if one had to collect data in multiple separate environments to vary a factor like table texture, but could collect data in a single environment to vary object type. In such situations, it may be easier to use the L strategy to collect data varying object type in only one environment, and collect data varying table texture in separate environments without varying object type, instead of varying both factors together like Stair and Diagonal would require.
III-D Hypotheses
We develop the following hypotheses, which we seek to investigate in our experiments:
  1. 1.
    Stair, L, and Diagonal may outperform Random by exploiting compositional generalization when possible.
  2. 2.
    Stair, L, and Diagonal may approach the performance of Complete when composition is strong.
  3. 3.
    Stair may outperform L and Diagonal, by capturing a greater quantity and diversity of factor combinations.
  4. 4.
    Incorporating prior robot data can promote stronger composition of factor values in the in-domain data.
IV Simulation Experiments
To evaluate composition and data collection strategies at a large scale, we first conduct extensive experiments in simulation. We use Factor World, a robotics simulation platform that supports varying different environmental factors [46].
Figure 3: Visualization of the Pick Place task. We show combinations of 2 values each for the factors table texture and object position.
Figure 4: Simulation results of data collection strategies for Pick Place. We compare strategies using the same amount of data for each. We evaluate every factor pair from 5 factors ($N=2$), average results across all pairs, and all 5 factors together ($N=5$). The strategies that exploit composition (Stair, L, Diagonal) generally outperform Random, and often approach the performance of Complete with enough factor changes. Stair generally performs the best, especially in the ($N=5$) setting. Error bars represent standard error across 5 seeds.
IV-A Evaluation Protocol
We consider Pick Place and Door Open, two commonly studied tasks in robotics. We consider 5 factors, including 4 that generalization was shown to be challenging for in the original Factor World experiments (object texture, table texture, camera position, and distractor objects), as well as object position, as an example of a factor that more strongly affects required physical motions. We show example factor value combinations for table texture and object position for Pick Place in Fig. 3, and examples of individual factors for both tasks in Appendix A. To investigate which specific factors can be composed by policies effectively, we study the setting where $\mathcal{F}^{N}$ consists of one of the 10 possible factor pairs ($N=2$). To investigate composition in a more realistic setting where robustness to multiple factors is desired, we also study when $\mathcal{F}^{N}$ consists of all factors ($N=5$). For each factor, we sample $k=10$ values that the factor can take in $\mathcal{F}^{N}$. For each data collection strategy, we use a scripted expert policy to collect multiple datasets with different total amounts of factor changes, by setting new factor values according to the strategy at different rates. For example, with the Random strategy, 20 factor changes would involve resampling new factor combinations roughly twice as often as with 10 factor changes. We compare how the strategies scale with the total number of factor changes, which we use to quantify effort.
We only compare against Complete in the $N=2$ setting, as $N=5$ would require a minimum of $10^{5}$ demonstrations, which is impractical. Each dataset consists of 100 demonstrations for the $N=2$ experiments, and 1000 demonstrations for $N=5$. We train a policy on each dataset using behavior cloning, and evaluate on 100 episodes, each with a different $f$ randomly sampled from $\mathcal{F}^{N}$, to assess overall robustness to factor combinations. For the $N=2$ experiments, this consists of all possible $f\in\mathcal{F}^{N}$. We evaluate across 5 random seeds, which include different possible factor values in $\mathcal{F}^{N}$ for each. We provide more details on these experiments in Appendix A.
IV-B Results
We illustrate our results for Pick Place in Fig. 4. We provide similar results for Door Open in Section A-A. We first focus on the $N=2$ setting. We see that No Variation (dashed gray) performs poorly overall, verifying that exposure to a variety of factor values is needed to generalize. The Single Factor strategies (shades of blue) also perform poorly for most factor pairs, verifying that most pairs require exposure to variation in both factors to generalize. Note that these strategies can vary only a single factor at most 10 times, leading to shorter lines in the plots than the others. We see that the strategies intended to exploit compositional generalization – Stair, L, Diagonal (shades of orange) – generally outperform Random (light purple) at all levels of factor changes. They also approach the performance of Complete (dashed green line) with only 20 factor changes (enough to capture all individual factor values), compared to the 100 changes needed for Complete. These results suggest that our policies exhibit strong pairwise compositional abilities for these factors, and that Stair, L, and Diagonal are able to exploit this for more efficient data collection. We also see that Stair does slightly better overall than L and Diagonal, possibly suggesting some benefit from a greater quantity and diversity of factor value combinations.
Figure 5: (left) Compositional success rate of different strategies.
(right) Generalization of strategies with increasing dataset sizes.
In the $N=5$ setting, we see that performance is lower overall, indicating that generalization is harder when more factors are being varied. Stair, L, and Diagonal still outperform Random, but now Stair is the only strategy that is able to substantially improve over it. This suggests that a greater quantity and diversity of factor value combinations is more important for composition when there are more factors.
Assessing Composition. In Fig. 5 (left), we more closely investigate composition for the strategies intended to exploit it (Stair, L, Diagonal), and how this scales with the number of factor values seen. Unlike before, here we report each policy’s compositional success rate: the success rate only on combinations of factor values where each factor value was seen individually, but not in the exact same combination. The previous results considered all possible combinations, to assess overall generalization. We report this metric averaged across all pairs from our $N=2$ evaluation. We see these strategies exhibit strong composition for all levels of factor changes, suggesting that composition does not require observing a large number of factor values. However, Stair does appear to induce composition slightly better than the other strategies.
Scaling with Dataset Size. In Fig. 5 (right), we investigate how generalization scales with dataset size. We focus on the $N=5$ setting, where generalization is more challenging, and therefore the impact of data scale is more pronounced. We report the success rate of policies trained on datasets from different strategies, which all have 50 factor changes (the amount needed for Stair, L, and Diagonal to capture all factor values), but different amounts of demonstrations. We also compare against No Variation, which does not scale well, verifying that simply more data is not enough to generalize. The other strategies do scale positively, but Stair does significantly better overall than the rest, and it is the only strategy that is able to significantly benefit when increasing from 500 to 1000 demonstrations. This suggests that having more data, even without greater factor diversity, can help facilitate improved compositional generalization, but having a greater quantity and diversity of factor combinations, which Stair provides, may be critical for enabling this scaling.
V Real Robot Experiments
Our simulation results have been encouraging regarding composition in robotic imitation learning, and the potential to exploit this with strategies for more efficient data collection. However, for such strategies to be useful for real-world data collection, real robot policies must also have compositional abilities. Therefore, we conduct experiments to evaluate if and when compositional generalization happens on a real robot. In addition, we include experiments incorporating prior robotic datasets, to investigate if access to such datasets can benefit composition in robotics, as is the case with natural language.
V-A Evaluation Protocol
Figure 6: Our WidowX robot setup. We consider the task of putting a fork inside a container, in a real office kitchen.
Robot Platform. We conduct experiments using a WidowX 250 6DOF robot arm. We adapt the hardware and control setup used in BridgeData V2 [44], except we directly mount the robot and over-the-shoulder RGB camera to a mobile table. While our hardware setup is not identical to the original, we still seek to leverage BridgeData V2 as prior data, which represents a realistic use case of a prior robotic dataset. To promote transfer, we tune our camera setup so that policies trained on only BridgeData V2 can sometimes perform basic pick-place tasks zero-shot, although we were unable to have it perform the exact tasks in our evaluation. We show our robot setup in Fig. 6, and provide more details in Section B-A.
Figure 7: Visualization of our primary real robot factors in BaseKitch. The top row shows our base factor values $f^{*}$. The other rows show all deviations from $f^{*}$ by one factor value. We highlight the portion of each image affected by the specified factor value.
Task and Factors. We consider the task of placing a fork inside a container, on the countertop of an office kitchen. We primarily consider the following factors: object (fork) type, object (fork) position, container type, table height, and table texture. Each factor has $k=4$ possible values. We visualize our base values $f^{*}$, and also every deviation from $f^{*}$ by one factor, in Fig. 7. We additionally consider the factors camera position and distractor objects in secondary experiments.
Compared to our simulation experiments, our real experiments have a greater emphasis on physical factors that affect required physical motions, as opposed to visual factors that a policy should be mostly invariant to. For instance, we only study object type, container type, and table height in real, which are physical factors that can affect required grasp and place motions. This is because it is more difficult to conduct large-scale simulation experiments for such factors, due to the challenge of automatically collecting demonstrations that account for them. We believe that addressing physical factors is especially important because they are likely more resistant to visual data augmentation, and thus may require more emphasis during data collection. Also, physical factors are less likely to benefit from using internet data for semantic reasoning [18, 6].
Experimental Setup. We adopt a different experimental setup for our real experiments, due to the much greater cost of collecting data, training policies, and evaluating. We collect datasets for only the $N=5$ setting with all of our primary factors. Because the objective of these experiments is to determine which factors can be composed by a real robot policy, we focus on data collection strategies intended to exploit composition. In particular, we evaluate Stair, because it performed the best in simulation, and L, because by collecting data for each factor separately, we can use this to study pairwise composition using a model trained on a single dataset. L may also be more practical in real-world situations where it is easier to collect data for factors separately. For these strategies, we collect 10 human demonstrations for each of the 16 combinations of factor values required to cover all values per factor, resulting in 160 total demonstrations for each strategy. We also compare against No Variation with the same number of demonstrations. We collect data exclusively in one kitchen, which we refer to as BaseKitch. We train policies either from scratch, or with the addition of BridgeData V2, to assess the impact of prior data on composition. We provide more details on our experimental setup in Section B-B.
Data Strategy L No Variation Train Method Bridge From Scratch Bridge Factor 1 Factor 2 Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Position 8/9 5/9 2/9 5/9 0/9 0/9 0/9 1/9 3/9 2/9 1/9 1/9 Object Type 8/9 8/9 8/9 5/9 5/9 4/9 4/9 3/9 2/9 Container Type 5/9 6/9 4/9 6/9 2/9 3/9 Table Height 4/9 3/9 1/9 Overall 59/90 28/90 22/90
TABLE I: Real robot pairwise composition results. When leveraging BridgeData V2 as prior data, a policy is able to compose the factor values present in the L data to succeed on 59/90 compositional combinations of factor values. Without prior data, the model is unable to compose nearly as effectively, with compositional success rate dropping by roughly half. Prior data alone is also not enough to generalize to these situations, as a policy trained with prior data on No Variation also performs poorly.
Training. We adapt the diffusion goal-conditioned behavior cloning implementation from the BridgeData V2 code. For policies that use BridgeData V2 as prior data, unless otherwise stated, we first pre-train a policy on only the prior data, and then co-fine-tune on a mixture of in-domain and prior data. We verify that all policies succeed with base factor values $f^{*}$ in BaseKitch. We provide more details in Section B-C.
V-B Pairwise Composition
In this evaluation, we study pairwise composition of factors in BaseKitch. We compare L and No Variation to evaluate composition. For each pair of factors, we evaluate on all compositional combinations of factor values for that pair with respect to the L dataset, and set the values for other factors to their base values in $f^{*}$. This results in 9 evaluation scenarios for each of the 10 pairs. [^2] We have 3 values per factor that are not in the base values $f^{*}$, resulting in $93^{2}=$ 9 combinations for each of the (${5\choose 2}=10$ pairs. For L, we compare both training from scratch, and using BridgeData V2 as prior data. For No Variation, we only evaluate training with BridgeData V2.
We report our results in Table I. We find that by training on L data and using BridgeData V2 as prior data, our policy is able to compose the factor values present in the L data to generalize to 59/90 of the compositional combinations of factor values. These results suggest that composition is possible in real-world robotic imitation learning, similar to our simulation results. However, unlike in simulation, leveraging prior data is critical for strengthening compositional abilities in real. Without prior data, the policy’s overall compositional success rate is roughly halved to 28/90, with success rates dropping for 9/10 pairs of factors. This resembles results in prior work that suggest end-to-end neural models can struggle with composition, but large pre-trained models can exhibit strong composition [51, 13]. We hypothesize that prior data may be needed in real, but not in simulation, due to additional minor factor variations inherent to real experiments, including for factors that we do not account for. Our greater emphasis on physical factors in real may also account for some of this discrepancy. However, prior data alone is not enough for generalization, as training on No Variation with prior data performs much worse than using L data, achieving a success rate of only 22/90. This suggests that with current prior robotic datasets, varied in-domain data is often still critical for generalization, motivating efficient in-domain data collection.
We note that composition is generally strongest for pairs where at least one factor is visual, as opposed to physical factors that affect required physical motion. We hypothesize this is because physical factors can interact in more complex ways, making it more challenging for policies to compose unseen combinations of them. For example, object position and table height have the weakest composition together, possibly because both significantly affect the required grasp motion, and thus composition requires executing completely unseen grasps. Similarly, composition for container type is the weakest for the value white cup (second row in Fig. 7), which is narrower and taller than the other containers, and therefore requires a different place motion. In particular, this composes poorly with object position and table height, both physical factors. In contrast, composition is generally the strongest for object type. Although object type can sometimes affect required grasp and place motions, especially when interacting with other physical factors, we hypothesize it composes well because the different forks we consider are similar enough to not require drastically different physical motions. Overall, these results suggest that even with current prior robotics datasets, composition between physical factors can still be challenging, and as a result more exhaustive coverage for these factors during in-domain data collection may be necessary.
Pre-trained Representations. Pre-trained visual representations have become popular for promoting generalization in robot learning. To assess their impact on composition, we additionally compare with using R3M, a representation intended for robotic manipulation, as a frozen visual encoder [33]. We evaluate on a subset of factor pairs from our complete pairwise evaluation. Specifically, we consider object type + table texture (the pair of factors from our complete evaluation that affect physical motion the least), object position + table texture (a pair where one factor is physical and the other is visual), and object position + table height (a pair where both factors are physical). In our results in Table II, we find that R3M performs poorly overall. Qualitatively, we observe that the trajectories are more jittery than when learning end-to-end, as also noted in prior work that tried R3M with diffusion-based policies [9]. This suggests that modern frozen visual representations are not effective for compositional generalization on real robots, and that end-to-end learning is superior for this.
Data Strategy L No Variation Train Method Bridge R3M From Scratch Bridge R3M Object Type + Table Tex 8/9 1/9 4/9 2/9 0/9 Object Pos + Table Tex 5/9 0/9 1/9 1/9 0/9 Object Pos + Table Height 2/9 1/9 0/9 1/9 0/9 Overall 15/27 2/27 5/27 4/27 0/27
TABLE II: Additional real robot pairwise composition results with comparison to R3M as a frozen visual representation.
Figure 8: Visualization of BaseKitch with base factor values $f^{*}$ (left), compared to our transfer conditions in CompKitch (top right) and TileKitch (bottom right), which inherently have different factor values (e.g., table texture) and other characteristics (e.g., distractor objects, lighting). We consider setups both with no further changes from $f^{*}$ beyond these inherent differences, as well as with additional changes.
Data Strategy Stair L No Variation Train Method Bridge (Co-FT) Bridge (FT) From Scratch Bridge (Co-FT) Bridge (FT) From Scratch Bridge (Co-FT) Bridge (FT) No Further Changes CompKitch 4/5 4/5 0/5 5/5 0/5 0/5 0/5 0/5 TileKitch 5/5 2/5 0/5 5/5 5/5 3/5 0/5 0/5 Fork Down + Blue Plate CompKitch 4/5 4/5 0/5 3/5 0/5 0/5 1/5 0/5 TileKitch 4/5 3/5 0/5 3/5 5/5 5/5 0/5 0/5 Wooden Fork + Pink Bowl CompKitch 3/5 5/5 0/5 2/5 0/5 0/5 0/5 0/5 TileKitch 4/5 3/5 0/5 3/5 5/5 2/5 0/5 0/5 Plastic Fork + Higher Table CompKitch 4/5 0/5 0/5 1/5 0/5 0/5 0/5 0/5 TileKitch 3/5 0/5 0/5 2/5 0/5 1/5 0/5 0/5 Overall CompKitch 15/20 13/20 0/20 11/20 0/20 0/20 1/20 0/20 TileKitch 16/20 8/20 0/20 13/20 15/20 11/20 0/20 0/20 Combined 31/40 21/40 0/40 24/40 15/40 11/40 1/40 0/40
TABLE III: Out-of-domain transfer results to new kitchens CompKitch and TileKitch. We find that using both varied in-domain data from BaseKitch and BridgeData V2 as prior data is critical for effective transfer to these new kitchens. Stair outperforms L, although both achieve significant levels of transfer. Co-fine-tuning generally performs better than only fine-tuning.
V-C Out-Of-Domain Transfer
To further demonstrate the utility of exploiting composition during data collection, we assess the transfer abilities of policies trained on our datasets to entirely new environments that capture some of the factor variety accounted for during data collection. We evaluate in two new kitchens, which we refer to as CompKitch and TileKitch. [^3] We name these kitchens based on the material of their countertops. See Fig. 13 in Section B-B for additional views that make this more apparent. These kitchens inherently have some completely out-of-distribution factor values from those studied in BaseKitch (e.g., table texture, shown in Fig. 8), including for some factors we do not account for during data collection (e.g., distractor objects, lighting). We first evaluate in these new kitchens with no further changes from the base factor values $f^{*}$ in BaseKitch, aside from their inherent shifts. In addition, to assess beyond pairwise composition for multiple factors, we also evaluate in these kitchens with additional combinations of factor shifts. We visualize these different conditions in Fig. 8. For each data collection strategy, we again compare using BridgeData V2 as prior data, and training from scratch. In addition to using prior data for both pre-training and co-fine-tuning (as done with the previously evaluated models), we also compare to only fine-tuning on in-domain data from a pre-trained BridgeData V2 policy, to assess the importance of co-fine-tuning for compositional generalization and transfer.
We report our results in Table III. We find that BridgeData V2 is critical for transfer, as 2 of the 3 policies trained from scratch are unable to transfer at all, while the other only achieves a success rate of 11/40. When co-fine-tuning with BridgeData V2, having varied in-domain data is still needed for robust transfer, as No Variation only achieves a success rate of 1/40. These results indicate that this transfer setting represents a significant and challenging domain shift.
Despite this, Stair and L both achieve significant levels of transfer, with success rates of 31/40 and 24/40, respectively. Stair generally outperforms L, as was the case in simulation. This suggests that policies that use prior data are able to effectively compose factor values seen in their training data to transfer to new settings that require composition. When fine-tuning from a pre-trained model, rather than co-fine-tuning, Stair and L produce policies that achieve some transfer, but less consistently than when co-fine-tuning, with reduced success rates of 21/40 and 15/40, respectively. In particular, the L policy fails to transfer to CompKitch, and both policies fail to transfer to Plastic Fork + Higher Table in both kitchens. This suggests that co-fine-tuning is generally a stronger method of leveraging prior data to facilitate composition than pre-training alone. Overall, we believe these results further suggest that policies can exhibit composition to generalize to unseen settings, our data collection strategies are sufficient to achieve some of this composition, and that using prior data is important for this composition to happen effectively.
V-D Additional Experiments
Unaccounted Factors. Our transfer experiments involve generalization to some factors that were unaccounted for during data collection, such as distractor objects and lighting. For more focused assessment of the impact of data collection strategies on such unaccounted factors, we additionally evaluate on distractor objects as a held-out factor in BaseKitch. In early experiments, we noticed a large degree of robustness to this factor from No Variation policies, so we did not include it among our primary factors. We evaluate our policies on 3 different values for this factor (consisting of distinct objects and positions for each value), which we visualize in Fig. 9. Our results in Table IV confirm the aforementioned robustness of the No Variation policies. However, when training from scratch, policies using Stair and L data perform worse, indicating that these data collection strategies may worsen robustness for unaccounted factors, possibly due to introducing spurious correlations. However, incorporating prior data completely mitigates this issue and restores robustness to this factor. This suggests another reason for why using prior robot data can be important for generalization: to address factors unaccounted for during in-domain data collection.
Data Strategy Stair L No Variation Train Method Bridge From Scratch Bridge From Scratch Bridge From Scratch Tape Measure 5/5 1/5 5/5 5/5 5/5 5/5 Pink Bowl 5/5 3/5 5/5 4/5 5/5 5/5 Spoon 5/5 5/5 5/5 0/5 5/5 4/5 Overall 15/15 9/15 15/15 9/15 15/15 15/15
TABLE IV: Evaluation on held-out factor distractor objects. Data collection strategies that do not account for this factor can perform worse than No Variation when training from scratch, but this issue is alleviated completely by using prior data.
(a) Tape Measure
(b) Pink Bowl
(c) Spoon
Figure 9: Visualization of the different values for distractor objects we consider in our held-out factor evaluation.
(a) Main Camera
(b) Secondary Camera
Figure 10: Visualization of the secondary camera view we use for studying composition of camera position (right), compared with our main camera view used in all other experiments (left).
Camera Position Composition. We do not consider camera position among the primary factors in our evaluation, due to the difficulty of varying this factor in a controlled manner. However, we have one additional third-person camera on our robot platform that we also collect data for, which allows us to separately study compositional generalization with this secondary camera. We visualize the difference in perspective with this secondary camera from our main camera in Fig. 10.
We focus on the composition of camera position with table texture. We evaluate in BaseKitch, but using the secondary camera for observations. We set table texture to the value white marble (third row in Fig. 7), as this was the value that was the most challenging for policies lacking variation for table texture to generalize to. We train a policy on the combination of two sub-datasets: our L dataset from the primary camera view, and the same L dataset but from the secondary camera view, with the exception of data for table texture. We compare against training on only one of these sub-datasets, which either lack data with the correct value for camera position, or the correct value for table texture, that is seen during evaluation. We again evaluate how using BridgeData V2 as prior data affects this composition. In our results in Table V, we find that yet again, policies that do not use prior data struggle to generalize. When using prior data, training on each sub-dataset individually results in a policy that sometimes generalizes, but not consistently. However, training on both sub-datasets together is able to achieve a perfect success rate, suggesting that the policy is able to effectively compose camera position with table texture.
Train Method Bridge From Scratch
No Camera Position 1/10 0/10
No Table Texture 6/10 0/10
Both 10/10 0/10
TABLE V: Our policy composes datasets missing either the correct camera position or table texture seen during evaluation, to outperform training on either dataset alone. Prior data is critical for composition.
VI Discussion
Summary. We investigate the compositional abilities of visual imitation learning policies for robotic manipulation, and whether this can be exploited by data collection strategies to more efficiently achieve broad generalization. In summary, our simulated and real-world experiments suggest the following:
  • Robot policies do exhibit significant composition, although this does not always happen on a real robot.
  • Leveraging prior robot data is critical for strengthening compositional abilities with a real robot, and co-fine-tuning is generally better for this than fine-tuning.
  • Our proposed data collection strategies are able to exploit composition to reduce data collection effort, while producing policies that can generalize to unseen settings.
  • Two of our proposed strategies, Stair and L, are able to transfer policies to entirely new environments that require a high degree of compositional generalization, showing they can produce data that is useful beyond the original domain it was collected in.
  • Our proposed Stair strategy is generally the most effective, as it achieves the best pairwise composition in simulation and real, the best transfer results in real, and is the only strategy in simulation that effectively scales with dataset size for the same amount of factor variation.
Overall, we believe these results provide insights on how roboticists can more efficiently collect in-domain data for achieving generalization in their desired settings.
Limitations and Future Work. While we conduct thorough real-world experiments for evaluating compositional generalization, it would be interesting to investigate how these results scale to more complex robot platforms and tasks. We did some preliminary investigation into treating different tasks as a factor for composition, by combining data for a new task “remove fork from container” with the L dataset for our main task “put fork in container”. However, we were unable to achieve significant composition in early experiments. This could be because our new task data did not have any factor variation at all, which might be necessary for composition. Future work can more thoroughly consider this possibility.
Although we believe our results can be broadly informative for data collection in robotics, we focus primarily on collecting in-domain data. Future work could scale this analysis to large-scale data collection efforts, to better understand how to most efficiently collect prior data for downstream transfer. While we do show that incorporating such prior data is critical for strengthening composition in our experiments, we only consider one prior dataset. Future work can investigate how other prior robotic datasets impact composition, and what aspects of prior datasets are the most important for this.
Finally, while our results show that robot policies can possess significant composition and transfer capabilities, they still struggle at times with composition, particularly for factors that interact physically. Future work can better address generalization to these settings, such as through data collection strategies that emphasize certain factor combinations over others. Also, we only consider straightforward behavior cloning methods for policy learning. It would be interesting to study if other approaches to policy learning have better compositional abilities, such as different learning algorithms or architectures.
Appendix A Simulation Experiments
A-A Door Open Results
Figure 11: Simulation results of different data collection strategies for Door Open. We include results for every pair from 5 factors ($N=2$), average results across all pairs, and all 5 factors together ($N=5$). Results are similar as with Pick Place, with Stair generally performing the best, especially in the ($N=5$) setting. Error bars represent standard error across 5 seeds.
We include additional simulation results for the task Door Open in Fig. 11. Results are similar to our results for Pick Place, with pairwise composition being strong for each pair of factors, and Stair generally performing the best, especially in the $N=5$ setting.
A-B Factors
For the inherently discrete-valued factors object texture, table texture, and distractor objects, we sample our $k$ discrete values from all possible training values specified in Factor World. Distractor objects also include a size scale (sampled from range $[0.3,0.8]$), 2D rotation (sampled from range $[0,2\pi]$), and 2D position (sampled from all possible positions on the table) as part of each value. For object position, we sample discrete 2D $xy$ positions from the range $[-0.1,0.1]$ for both coordinates. We note that the Pick Place task includes a small amount of added noise to object position each episode, sampled uniformly from the range $[-0.03,0.03]$. For camera position, we sample $xyz$ positions and rotation quaternions all from the range $[-0.05,0.05]$. When sampling our $k$ discrete values for each factor, we ensure that the scripted policy is able to solve the task for each value, because some values for object position and distractor objects can impede the task. In Fig. 12, we visualize two examples of values for each factor, for both the Pick Place and Door Open tasks. We note that each random seed in our evaluation includes a different set of $k$ discrete values sampled for each factor.
A-C Training
We use the same policy architecture and training hyperparameters used in the original Factor World experiments [46]. We condition policies on the same observations (2 84x84 RGB image observations from 2 different camera views without history, and proprioception). Unlike in the original Factor World experiments, we use random shift augmentations in all experiments. Unlike our real robot experiments, we train policies without goal conditioning, as we do not experiment with incorporating diverse prior data in this setting, which would benefit from some form of task conditioning.
A-D Computing Factor Changes
For our results in Section IV-B, we compute the total number of factor changes for each strategy as follows. We assume the initial configuration of factor values requires $n$ changes, one for each factor. For Stair, L, and Single Factor, we assume each new configuration of factor values requires 1 additional change. We note this could be a slight underestimate for L in practice, as this strategy could require some additional changes when finishing varying one factor and returning to the base factor values $f^{*}$, to begin data collection for varying another factor. For Diagonal and Random, we assume each new configuration of factor values requires $n$ additional changes, as new values for all factors are resampled. We note that this could be a slight overestimate for Random in practice, as some of the resampled factor values may not change.
For each budget of total number of factor value changes, we determine the amount of factor value configurations allowed for each strategy. We then divide the budget of total demonstrations by this to determine how many demonstrations to collect for each configuration, with any remainder going to the last configuration.
(a) Object Position
(b) Object Texture
(c) Table Texture
(d) Camera Position
(e) Distractor Objects
Figure 12: Visualization of our factors for the Pick Place (left) and Door Open (right) tasks from Factor World. We show two examples of values for each factor we consider.
Appendix B Real Robot Experiments
B-A Robot Platform
We use Logitech C920 webcams for our main and secondary third-person cameras. Our setup also includes the wrist camera used in BridgeData V2, and we experimented with using it in addition to our main third-person camera. While we found that using it improved overall robustness when training from scratch, particularly for some factors that wrist cameras provide invariance to (e.g., object position, table height), we also found that it was incompatible with using BridgeData V2 as prior data, reducing performance compared to not using prior data at all. We believe this is because only a small fraction of this dataset contains wrist camera data, as similar negative results with wrist cameras have been observed in other work that use prior robotic datasets with a small proportion of wrist camera data [34]. As we were only able to achieve successful transfer in our experiments in Section V-C by using prior data without a wrist camera, we decided to omit the wrist camera in all our experiments.
(a) BaseKitch
(b) CompKitch
(c) TileKitch
Figure 13: Additional views of each kitchen in our real robot experiments. We name CompKitch and TileKitch after their countertop materials of composite and tile, respectively. These images were taken after our evaluations, so there may be some slight differences from then.
B-B Evaluation Protocol
For our out-of-domain transfer experiments in Section V-C, we name our transfer kitchens CompKitch and TileKitch after their countertop materials of composite and tile, respectively. We provide additional views of these kitchens, as well as BaseKitch, in Fig. 13 to make the difference between these kitchens more apparent. These images were taken after our evaluations, so there may be some slight differences from then.
We vary the factor table height by adjusting the height of the mobile table the robot and third-person cameras are mounted on. In other words, this factor is varied in practice by varying the height of the robot’s table, not the object table (the surface that objects are being manipulated on). However, this still changes the relative height of the robot with respect to the object table, the same as if the height of the object table was adjusted instead. In Section V-C, Higher Table refers to the object table being higher relative to the robot, which was achieved by lowering the robot table.
We collect demonstrations using a Meta Quest 2 VR headset for teleoperation. All demonstrations are collected by a single experienced human teleoperator for consistency. We collect our Stair dataset by starting at $f^{*}$, and then varying factors cyclically in the order object position, object type, container type, table height, table texture. The order in which we vary values for each factor is the same as in Fig. 7, from top to bottom.
B-C Training
We use the same ResNet-34 diffusion goal-conditioned policy architecture used in the original BridgeData V2 experiments, except we condition on a history of 2 128x128 RGB image observations from a third-person view (in addition to a goal image from the same view), and use action chunking to predict the next 4 actions. However, during inference we noticed better performance by only executing the first predicted action, so we do this for all experiments unless otherwise stated. We share the same visual encoder across image observations in the history. We use the same data augmentation procedure used in the original BridgeData V2 experiments, which consists of random crops, random resizing, and color jitter.
For policies using BridgeData V2 as prior data, we first pre-train a model on prior data using the original training hyperparameters for 2M gradient steps, and choose the checkpoint with the lowest validation action prediction mean-squared error as the initialization for later fine-tuning. During fine-tuning, we use the same pre-training hyperparameters, except we train for only 300K gradient steps and use the final checkpoint. When co-fine-tuning, we train on a mixture of 75% in-domain data and 25% prior data.
B-D R3M
We use the ResNet-50 version of R3M. When training, we pre-compute representations beforehand, and standardize the dataset such that each feature has mean 0 and standard deviation 1 (similar to the batch normalization used in the original R3M experiments). We use the training dataset mean and standard deviation for normalization during inference. We feed normalized representations to the same diffusion policy head architecture used when learning end-to-end, except we do not use goal-conditioning. We increase the amount of training gradient steps from 300K to 500K, to reduce training loss. We also tried 1M gradient steps, which reduces loss even further, but this resulted in worse performance. Instead of executing only the first predicted action as with the end-to-end policies, we execute all 4, which we found to slightly reduce jitteriness. Like when learning end-to-end, we verify our R3M policies are able to succeed with base factor values $f^{*}$ in BaseKitch.